Jefferson Lab Weekly Briefs April 4, 2012
Jefferson Lab Weekly Briefs
April 4, 2012
12 GeV Upgrade
Installation of the new coldboxes for Central Helium Liquefier-2 has begun by the vendor representative. The C100-1 and C100-2 cryomodules continue to be used for beam operations; refinement of the control of the radiofrequency fields is making good progress. Testing of the fifth new cryomodule is underway in the Cryomodule Test Facility. The procurement process for the re-bid of the box power supplies (for the large dipoles in the beam transport system) is underway. While this will delay delivery relative to the earlier plan, we do not presently project any impact on the commissioning schedule. Manufacturing of the new stripline-design beam position monitor cans is underway.
Hall A recovered from the septum incident and locked up the hall on Tuesday to accept beam delivery. After an initial period of optics calibrations and tests of low-current beam instrumentation, the hall began production data-taking on Thursday for the g2p experiment. The hall ran smoothly all weekend.
In Hall C, the Q-weak experiment continued data taking on liquid hydrogen with the 180 microamp beam. Maintenance activities included replacing the power supply for the Moller polarimeter's superconducting solenoid and repairing a slow water leak internal to the Qtor power supply.
Last Chance to Submit Nominations for Assistantship
The week started with a maintenance day, which included a search for the reason for bad optics in the five-pass setup. A faulty control card for a focusing quadrupole was determined to be the source and was replaced. On Thursday, Hall B was successfully set up for five-pass beam. Filament power supply boards caused some problems with the radiofrequency amplifiers. Double peaking in Hall C data prompted several attempts to adjust the timing of the accelerating cavities to correct. Orbit drifting in the injector caused the vacuum valves to shut. The orbit was corrected and beam ran fine into the weekend. The week ended with a third fire alarm, this time in Hall C, that was due to mist coming out of a de-humidifier machine. Halls A, B and C took beam for several days in a row without any major incidents, then while Hall B was working on their target, Halls A and C continued receiving continuous wave beam on varying targets. The gun cathode was re-activated, and new injector serial communications were installed for the injector beam current monitors.
The FEL team continued with extensive beam operations last week, performing a number of electron beam and optics measurements for the Office of Naval Research. Extended second shift beam operations successfully delivered terahertz and infrared beam to outside users looking at insulator-metal transition and enhanced molecular resonance absorption. Engineering work on the new gun continues, as does engineering for upgrades of the cryomodule to be installed in zone FL03 late next fall.
Center for Theoretical and Computational Physics
Extracting information about hadron resonances from lattice QCD calculations involves measurement of the discrete spectrum of states in a finite-size box. As a first step toward this, Theory Center members have extracted the non-resonant elastic scattering phase shift for two pions with isospin 2 in both S- and D-waves. A detailed determination of the energy dependence follows from the use of a number of novel lattice field theory techniques, and the result of a weak repulsive interaction is in line with experimental data.
The Machine Shop has processed 166 job requests since Feb. 20. It is presently manufacturing or continuing to manufacture:
Facilities Management and Logistics
JLab Gets New Street, Building and Informational Signs
JLab Calendar of Events
April 5: Seminar: Women in Physics
Environment, Safety, Health and Quality
Resolving ES&H Concerns
Once your concern has been addressed, communicate any lessons learned to your staff, co-workers, supervisors and your Division Lessons Learned Coordinator. ES&H Manual Chapter 2310 offers more information on concern resolution.
Heavy Traffic at CEBAF Center Circle Midday on April 5
Tornado Warning Siren Test Set for Noon Friday
JLab in the News
Seminar on Women in Physics Offered April 5
Volunteers Needed for May 19 JLab Open House
Tracking Thomas at JLab
Congratulations this week go to Maurizio Ungaro and Samika Hawkins, who were the first to correctly identify the March 28 location. Honorable mentions go to Doug Higinbotham, Michael Haddox-Schatz, Dan Moore, Rob Mahurin, Anusha Liyanage, Andrea Celentano, Lori Zukerman, Annie Tran, Brian Agor and Subashini De Silva. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.